# What Changed and What Could Have Changed? State-Change Counterfactuals for Procedure-Aware Video Representation Learning

### v3

@@ Abstract
Understanding a procedural activity requires modeling both how action steps transform the scene, and how evolving scene transformations can influence the sequence of action steps, even those that are accidental or erroneous. Existing work has studied procedure-aware video representations by proposing novel approaches such as modeling the temporal order of actions and has not explicitly learned the state changes (scene transformations). In this work, we study procedure-aware video representation learning by incorporating state-change descriptions generated by Large Language Models (LLMs) as supervision signals for video encoders. Moreover, we generate state-change counterfactuals that simulate hypothesized failure outcomes, allowing models to learn by imagining the unseen “What if” scenarios. This counterfactual reasoning facilitates the model’s ability to understand the cause and effect of each step in an activity. To verify the procedure awareness of our model, we conduct extensive experiments on procedure-aware tasks, including temporal action segmentation, error detection, action phase classification, frame retrieval, multi-instance retrieval, and action recognition. Our results demonstrate the effectiveness of the proposed state-change descriptions and their counterfactuals, and achieve significant improvements on multiple tasks. We will make our source code and data publicly available soon.
@@ 1 Introduction
Procedural activities, such as following a cooking recipe [gtea, breakfast, egoper, grauman2024ego] or assembling a piece of furniture [miech2019howto100m, sener2022assembly101], consist of sequences of interconnected steps that often happen in a specific and logical order to achieve a desired outcome. Understanding these activities from video data is essential for a variety of applications, including video retrieval [dong2019dual, miech2019howto100m, gabeur2020multi, Wray_2021_CVPR], intelligent collaborative agents [kung2023riskbench, wang2024mosaic, banerjee2024ask, ye2024morpheus, jenamani2024feel], and robot learning from human demonstration [7102751, shao2021concept2robot, xiao2022masked, zakka2022xirl, xu2023xskill, zhu2024orion, jain2024vid2robot, ren2025motion]. In contrast to general video action recognition that focuses on a single step in a short clip [carreira2017quo, feichtenhofer2019slowfast, fan2021multiscale, wang2021actionclip, tong2022videomae, Kung_2024_CVPR], procedure-aware video understanding requires capturing both the “what changed” (actual action-induced state transformations) [liang2022var, souvcek2023genhowto], as shown in Figure 1 (a), and “what could have changed” (hypothetical deviations) [HARRIS1996233, roese1997counterfactual, schulam2017reliable, zhang2021if], as shown in Figures 1 (b) and (c). These two capabilities play crucial roles in understanding long-form procedures where each action step can have a sequential effect on the others.
@@
Procedure-aware video representation learning can enable various downstream long-form video tasks, such as temporal action segmentation [gtea, breakfast, sener2022assembly101, egoper] and error detection [sener2022assembly101, egoper]. There have been many approaches proposed to learn procedure-aware representations, including learning spatiotemporal features [carreira2017quo, feichtenhofer2019slowfast, fan2021multiscale, tong2022videomae], using action labels as supervision [miech2020end, lin2022learning, xiao2022hierarchical, zhong2023learning], incorporating temporal order of steps [jenni2020video, zhong2023learning], and consulting external activity procedure knowledge databases [zhou2023procedure] such as WikiHow [koupaee2018wikihow]. However, these approaches often struggle to explicitly model procedural knowledge that relies on how scene states evolve over time and how they could have evolved if the precondition or result of an action step had differed. For instance, the procedure of “Make Oatmeal” in Figure 1 (c) cannot be finished correctly if the state of the bowl remains “Full of oatmeal and without water” or “Full of water without oatmeal” because they do not meet the precondition of the later action step, “Heat with microwave”. Capturing these subtle transitions requires state-change reasoning to understand how an action transforms an environment or objects, and counterfactual analysis to foresee alternative outcomes for the same key action.
@@
In this paper, we propose a hierarchical procedure-aware video representation learning framework that leverages state-change and their counterfactuals generated by a Large Language Model (LLM) [dubey2024llama]. We use the actual state-change descriptions to temporally contrast features, and the counterfactuals as the negative exemplars in contrastive learning [oord2018representation, chen2020simple, 10.5555/3495724.3497291] to enhance the modeling of clip-level and video-level procedure transformations. Specifically, at the clip level, we first capture before and after state that describe the change of environments or objects to identify local action-induced state changes. Moreover, we also incorporate state-change counterfactuals that estimate a hypothesized state if the action step were to fail for any reason. For example, if the action were “make coffee,” a possible state counterfactual could be “cup shards scattered on the floor.” We then use temporal contrastive learning to encourage the model to learn a representation space in which visual features of later frames in short clips and after states are close together, and in which the early frames, before state, and state counterfactual are far away, and vice versa. At the video level, we further incorporate counterfactual reasoning by generating missing-step and misordered, enabling a more comprehensive understanding of the entire procedure in a long-form video.
@@
We evaluate the learned procedure-aware video representations in six key procedural and short-term video understanding tasks—error detection, temporal action segmentation, multi-instance retrieval, action recognition, action phase classification, and frame retrieval—and show that they significantly enhance performance compared to strong baselines. To validate our approach, we incorporate state changes and their counterfactuals into the learning process and pre-train our model on Ego4D [grauman2022ego4d, kevin2022egovlp], a large-scale daily activity dataset filtered to include clean and aligned short clips along with their corresponding clip-level narration text and video-level summary text. We then evaluate our model on various procedure-aware downstream tasks, including long-form video tasks like temporal action segmentation and error detection, and short-term action recognition and retrieval. In addition, we conduct a comprehensive ablation study to highlight the impact of learning state changes and their counterfactuals, demonstrating their importance in enhancing procedural video understanding. Our contributions are summarized as follows:
@@ 2.1 Video Procedure-Aware Representations
Procedure understanding, which involves recognizing and localizing individual action steps and reasoning about their causal relationships, is a fundamental challenge in computer vision, and is important for applications such as step recognition [tang2019coin, damen2020epic, bansal2022my], temporal action segmentation [gtea, breakfast, sener2022assembly101, chen2025atars], and error detection [egoper]. Early work proposed learning procedure-aware video representations via exploiting the temporal consistency within videos [wang2015unsupervised, bansal2022my, xiao2022hierarchical, zhang2022unsupervised, qing2022learning] and temporal order [jenni2020video] as pre-text tasks. Some work incorporates task graphs as guidance to learn procedure-aware representations, either by constructing graphs based on knowledge bases [zhou2023procedure] or through learning [narasimhan2023learning, mavroudi2023learning, bansal2024united, seminara2024differentiable, ashutosh2024video]. Recently, several papers have attempted to align video features in datasets such as HowTo100M [miech2019howto100m] or Ego4D [grauman2022ego4d] with text descriptions extracted through automatic speech recognition [miech2020end, mavroudi2023learning], refined subtitles [lin2022learning] with the external database WikiHow [koupaee2018wikihow, mavroudi2023learning], or manually labeled annotations [kevin2022egovlp, zhong2023learning, ashutosh2023hiervl]. These text descriptions offer rich context of action steps and temporal relations, opening the door for procedure-awareness learning. For example, Br-prompt [li2022bridge] generates text prompts from ground truth action labels to learn ordinal temporal features. PVRL [zhong2023learning] formulates future action steps as next-token prediction [devlin2019bert], enhancing temporal order between actions. HierVL [ashutosh2023hiervl] proposes hierarchical video-text alignment for both short clips and long videos.
However, these papers fall short of explicitly learning state changes—the causal relationships between action and state—which is essential to procedure understanding [souvcek2023genhowto, niu2024schema]. Moreover, they overfit the seen correctly executed actions, which hinders the understanding of activity procedures and the ability to recognize the deviation of actions, such as erroneous action steps. In this work, we propose to incorporate state-change descriptions and counterfactuals to force our model to learn about action transformations and hypothetical outcomes that facilitate “learning from hypothetical failures” for procedure awareness.
@@ 2.2 State-Change Understanding
The idea of visual state changes has been explored extensively in the computer vision community, sometimes using different terms such as transformations between states [wang2016actions, liang2022var] or state-modifying actions [souvcek2022multi, souvcek2022look, niu2024schema, soucek2024multitask]. For example, papers have studied object state change localization [liu2017jointly, alayrac2017joint, souvcek2022look, saini2023chop, xue2024learning] and generation [soucek2024multitask], and action recognition with state changes [fathi2013modeling, wang2016actions, saini2023chop]. Procedure planning, which aims to generate effective plans for task completion based on start and goal visual states, is another task that heavily incorporates the concept of understanding state changes [chang2020procedure, lu2022neuro, wang2023event, li2023skip, niu2024schema, nagasinghe2024not]. We were also inspired by SCHEMA [niu2024schema] to use LLMs to generate before- and after-state descriptions of actions. However, three key factors sets our work apart from the existing work. First, unlike existing work that mainly focuses on a specific task such as procedure planning [niu2024schema], we aim to design a more general video representation learning framework for various downstream tasks. Second, prior work does not incorporate hypothetical outcomes, i.e., counterfactuals, which are essential for procedure causal reasoning. Last but not least, prior work often only studies interacted objects’ state changes. In contrast, our generated descriptions can cover state changes of objects, humans, and the surrounding environment. We provide extensive examples in our supplementary materials.
@@ 2.3 Video Representations with Counterfactuals
Counterfactual example reasoning that helps models understand cause-and-effect and predict outcomes [scholkopf2021toward, yang2021causalvae, zang2023discovering] has demonstrated strong generalization and robustness on various tasks including video question-answering [CLEVRER, li2022representation, zhang2024if] and compositional reasoning [sun2021counterfactual, lai2024improving]. Extensive research has proposed to enhance the causality in visual learning by creating counterfactual examples with image data augmentation [zhong2020random], text counterfactuals based on rule-based [shi2018learning], random word substitution [roth2023waffling], language models [doveh2023teaching, wang2023vilta, xie2023ra, zhang2024countercurate], and synthetic image-text pairs generated by a simulator [cascante2023going] or generative model [shi2023logoprompt, le2023coco, lai2024improving], such as stable diffusion [rombach2022high]. However, existing work mainly focuses on creating counterfactuals for images or short clips and does not consider the procedural nature of activities. In this work, we propose state-change counterfactuals at both the clip and video levels, highlighting the action transformations and procedure evolution.
@@ 3.1 State Change & Counterfactual Generation
We use the LLM Llama 3 [dubey2024llama] to generate clip-level state descriptions, i.e., before state, after state and state-change counterfactual, and to generate video-level summary counterfactuals, i.e., missing step and misordered steps. In the following, we introduce the prompts for text generation.
Clip-Level State Changes and their Counterfactuals. We feed the narration annotated for each short clip in Ego4D [grauman2022ego4d, kevin2022egovlp] into Llama and ask it to generate a state-change description based on the given narration. Specifically,
@@
Video-Level Counterfactual Generation. We feed all clip narrations in a long video and their video-level summaries annotated in Ego4D [grauman2022ego4d, kevin2022egovlp] to Llama. We ask the LLM to generate summary counterfactuals for missing key steps [K-CF] in which some crucial action steps (clip narrations) are missing,
@@ 3.2 Preliminary: Hierarchical Video
In this work, we adopt HierVL [ashutosh2023hiervl], a framework that learns hierarchical video-language representations at two temporal scales, clip-level and video-level. The clip-level narrations and video-level summaries are annotated by Ego4D/Egoclip [grauman2022ego4d, kevin2022egovlp]. Note that since HierVL trains both a visual and text encoder, it requires vision-to-text and text-to-vision alignment. Below we describe the former, and the latter is defined symmetrically. Formally, let a video be divided into $M$ short clips {$v_{1},\ldots,v_{M}$} with corresponding narration texts {$t_{1},\ldots,t_{M}$}, and let $S$ be the video-level summary for the entire video. HierVL’s vision-to-text training objective consists of two parts: clip-level alignment and video-level alignment losses.
@@
where $B$ is the mini-batch, $v_{i}$ and $t_{j}$ are visual and textual embeddings, respectively, and $P(i)$ denotes the positive samples of the $i^{\text{th}}$ video clip. HierVL [ashutosh2023hiervl] leverages a variant of the action-aware loss, EgoNCE [kevin2022egovlp], in order to mine hard positives and negatives; please refer to [ashutosh2023hiervl, kevin2022egovlp] for more details about this sampling. The overall loss at the child level is thus $\mathcal{L}_{\text{child}}=\mathcal{L}_{{v2t}}+\mathcal{L}_{{t2v}}$, where the latter is defined symmetrically.
@@
Br-prompt [li2022bridge] GTEA 94.1 92.0 83.0 91.6 81.2 88.4 - - - -
I3D [carreira2017quo] Kinetics 90.1 88.8 79.2 84.6 79.7 84.5 48.8 71.9 73.9 64.9
CLIP [radford2021learning] WIT [srinivasan2021wit]+Text encoder 88.5 86.2 77.6 87.1 75.6 83.0 44.2 71.2 70.8 62.1
MIL-NCE [miech2020end] HowTo100M 67.9 61.3 44.6 67.9 58.3 60.0 47.3 69.1 73.6 63.3
PVRL [zhong2023learning] HowTo100M 85.2 82.6 72.2 81.1 71.2 78.5 45.6 73.2 73.4 64.1
HierVL [ashutosh2023hiervl] Ego4D+Text encoder 90.4 88.5 81.2 86.7 78.5 85.1 52.6 73.0 77.3 67.6
@@ 3.3 Pretraining Objective:
Frame-Level Alignment. At the frame level, the model is supervised by our proposed before-state loss $\mathcal{L}_{\text{before}}$ and after-state loss $\mathcal{L}_{\text{after}}$, enhancing action-induced transformation for procedural understanding. The mathematical formulation of each is [10.5555/3495724.3497291],
@@
Clip-Level Alignment. At the clip-level, we leverage the $L_{v2t}$ loss described in Section 3.2, which seeks to align the video-clip embeddings to their corresponding text narrations from EgoClip [kevin2022egovlp]. Note that since we do not train a text encoder, the symmetric $L_{t2v}$ loss is neglected here. For more details on this loss, see Section 3.2 and [ashutosh2023hiervl, kevin2022egovlp]. The resulting loss for the first two scales is therefore
@@
where $V_{i}$ is the aggregated video-level visual embedding, $S_{j}$ is a summary text embedding and $S_{j,w}^{cf}$ are Misordered and Missing-step counterfactual text embeddings where $w\in\{1,\ldots,W\}$ is the total number of counterfactuals used, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ video, respectively. We also use a temperature hyperparameter $\tau$ as in Eq. (3), but we omit it in the equation above for simplicity of notation. Note that consistent with [ashutosh2023hiervl], in this case, the summation over positive samples is located inside the logarithm since the sampling strategy of [kevin2022egovlp] is used. However, prior work [10.5555/3495724.3497291] finds the summation outside the logarithm to be more effective and thus this is the de-facto choice of the frame-level loss in Eq. (3).

### v4

@@ Abstract
Understanding a procedural activity requires modeling both how action steps transform the scene, and how evolving scene transformations can influence the sequence of action steps, even those that are accidental or erroneous. Existing work has studied procedure-aware video representations by proposing novel approaches such as modeling the temporal order of actions and has not explicitly learned the state changes (scene transformations). In this work, we study procedure-aware video representation learning by incorporating state-change descriptions generated by Large Language Models (LLMs) as supervision signals for video encoders. Moreover, we generate state-change counterfactuals that simulate hypothesized failure outcomes, allowing models to learn by imagining the unseen “What if” scenarios. This counterfactual reasoning facilitates the model’s ability to understand the cause and effect of each step in an activity. To verify the procedure awareness of our model, we conduct extensive experiments on procedure-aware tasks, including temporal action segmentation, error detection, action phase classification, frame retrieval, multi-instance retrieval, and action recognition. Our results demonstrate the effectiveness of the proposed state-change descriptions and their counterfactuals, and achieve significant improvements on multiple tasks. Code is available at https://github.com/HCIS-Lab/counterfactual-video-pretrain.
@@ 1 Introduction
Procedural activities, such as following a cooking recipe [20, 32, 37, 24] or assembling a piece of furniture [49, 65], consist of sequences of interconnected steps that often happen in a specific and logical order to achieve a desired outcome. Understanding these activities from video data is essential for a variety of applications, including video retrieval [15, 49, 22, 85], intelligent collaborative agents [33, 80, 4, 94, 27], and robot learning from human demonstration [30, 66, 88, 97, 90, 106, 26, 57]. In contrast to general video action recognition that focuses on a single step in a short clip [7, 21, 18, 81, 78, 34], procedure-aware video understanding requires capturing both the “what changed” (actual action-induced state transformations) [41, 73], as shown in Figure 1 (a), and “what could have changed” (hypothetical deviations) [25, 58, 63, 102], as shown in Figures 1 (b) and (c). These two capabilities play crucial roles in understanding long-form procedures where each action step can have a sequential effect on the others.
@@
Procedure-aware video representation learning can enable various downstream long-form video tasks, such as temporal action segmentation [20, 32, 65, 37] and error detection [65, 37]. There have been many approaches proposed to learn procedure-aware representations, including learning spatiotemporal features [7, 21, 18, 78], using action labels as supervision [50, 43, 87, 103], incorporating temporal order of steps [28, 103], and consulting external activity procedure knowledge databases [105] such as WikiHow [31]. However, these approaches often struggle to explicitly model procedural knowledge that relies on how scene states evolve over time and how they could have evolved if the precondition or result of an action step had differed. For instance, the procedure of “Make Oatmeal” in Figure 1 (c) cannot be finished correctly if the state of the bowl remains “Full of oatmeal and without water” or “Full of water without oatmeal” because they do not meet the precondition of the later action step, “Heat with microwave”. Capturing these subtle transitions requires state-change reasoning to understand how an action transforms an environment or objects, and counterfactual analysis to foresee alternative outcomes for the same key action.
@@
In this paper, we propose a hierarchical procedure-aware video representation learning framework that leverages state-change and their counterfactuals generated by a Large Language Model (LLM) [17]. We use the actual state-change descriptions to temporally contrast features, and the counterfactuals as the negative exemplars in contrastive learning [54, 10, 29] to enhance the modeling of clip-level and video-level procedure transformations. Specifically, at the clip level, we first capture before and after state that describe the change of environments or objects to identify local action-induced state changes. Moreover, we also incorporate state-change counterfactuals that estimate a hypothesized state if the action step were to fail for any reason. For example, if the action were “make coffee,” a possible state counterfactual could be “cup shards scattered on the floor.” We then use temporal contrastive learning to encourage the model to learn a representation space in which visual features of later frames in short clips and after states are close together, and in which the early frames, before state, and state counterfactual are far away, and vice versa. At the video level, we further incorporate counterfactual reasoning by generating missing-step and misordered, enabling a more comprehensive understanding of the entire procedure in a long-form video.
@@
We evaluate the learned procedure-aware video representations in six key procedural and short-term video understanding tasks—error detection, temporal action segmentation, multi-instance retrieval, action recognition, action phase classification, and frame retrieval—and show that they significantly enhance performance compared to strong baselines. To validate our approach, we incorporate state changes and their counterfactuals into the learning process and pre-train our model on Ego4D [23, 42], a large-scale daily activity dataset filtered to include clean and aligned short clips along with their corresponding clip-level narration text and video-level summary text. We then evaluate our model on various procedure-aware downstream tasks, including long-form video tasks like temporal action segmentation and error detection, and short-term action recognition and retrieval. In addition, we conduct a comprehensive ablation study to highlight the impact of learning state changes and their counterfactuals, demonstrating their importance in enhancing procedural video understanding. Our contributions are summarized as follows:
@@ 2.1 Video Procedure-Aware Representations
Procedure understanding, which involves recognizing and localizing individual action steps and reasoning about their causal relationships, is a fundamental challenge in computer vision, and is important for applications such as step recognition [77, 13, 5], temporal action segmentation [20, 32, 65, 11], and error detection [37]. Early work proposed learning procedure-aware video representations via exploiting the temporal consistency within videos [83, 5, 87, 99, 55] and temporal order [28] as pre-text tasks. Some work incorporates task graphs as guidance to learn procedure-aware representations, either by constructing graphs based on knowledge bases [105] or through learning [52, 48, 6, 64, 3]. Recently, several papers have attempted to align video features in datasets such as HowTo100M [49] or Ego4D [23] with text descriptions extracted through automatic speech recognition [50, 48], refined subtitles [43] with the external database WikiHow [31, 48], or manually labeled annotations [42, 103, 2]. These text descriptions offer rich context of action steps and temporal relations, opening the door for procedure-awareness learning. For example, Br-prompt [39] generates text prompts from ground truth action labels to learn ordinal temporal features. PVRL [103] formulates future action steps as next-token prediction [14], enhancing temporal order between actions. HierVL [2] proposes hierarchical video-text alignment for both short clips and long videos.
However, these papers fall short of explicitly learning state changes—the causal relationships between action and state—which is essential to procedure understanding [73, 53]. Moreover, they overfit the seen correctly executed actions, which hinders the understanding of activity procedures and the ability to recognize the deviation of actions, such as erroneous action steps. In this work, we propose to incorporate state-change descriptions and counterfactuals to force our model to learn about action transformations and hypothetical outcomes that facilitate “learning from hypothetical failures” for procedure awareness.
@@ 2.2 State-Change Understanding
The idea of visual state changes has been explored extensively in the computer vision community, sometimes using different terms such as transformations between states [84, 41] or state-modifying actions [72, 71, 53, 74]. For example, papers have studied object state change localization [45, 1, 71, 61, 92] and generation [74], and action recognition with state changes [19, 84, 61]. Procedure planning, which aims to generate effective plans for task completion based on start and goal visual states, is another task that heavily incorporates the concept of understanding state changes [9, 47, 79, 40, 53, 51]. We were also inspired by SCHEMA [53] to use LLMs to generate before- and after-state descriptions of actions. However, three key factors sets our work apart from the existing work. First, unlike existing work that mainly focuses on a specific task such as procedure planning [53], we aim to design a more general video representation learning framework for various downstream tasks. Second, prior work does not incorporate hypothetical outcomes, i.e., counterfactuals, which are essential for procedure causal reasoning. Last but not least, prior work often only studies interacted objects’ state changes. In contrast, our generated descriptions can cover state changes of objects, humans, and the surrounding environment. We provide extensive examples in our supplementary materials.
@@ 2.3 Video Representations with Counterfactuals
Counterfactual example reasoning that helps models understand cause-and-effect and predict outcomes [62, 93, 98] has demonstrated strong generalization and robustness on various tasks including video question-answering [96, 38, 101] and compositional reasoning [76, 35]. Extensive research has proposed to enhance the causality in visual learning by creating counterfactual examples with image data augmentation [104], text counterfactuals based on rule-based [68], random word substitution [60], language models [16, 82, 89, 100], and synthetic image-text pairs generated by a simulator [8] or generative model [67, 36, 35], such as stable diffusion [59]. However, existing work mainly focuses on creating counterfactuals for images or short clips and does not consider the procedural nature of activities. In this work, we propose state-change counterfactuals at both the clip and video levels, highlighting the action transformations and procedure evolution.
@@ 3.1 State Change & Counterfactual Generation
We use the LLM Llama 3 [17] to generate clip-level state descriptions, i.e., before state, after state and state-change counterfactual, and to generate video-level summary counterfactuals, i.e., missing step and misordered steps. In the following, we introduce the prompts for text generation.
Clip-Level State Changes and their Counterfactuals. We feed the narration annotated for each short clip in Ego4D [23, 42] into Llama and ask it to generate a state-change description based on the given narration. Specifically,
@@
Video-Level Counterfactual Generation. We feed all clip narrations in a long video and their video-level summaries annotated in Ego4D [23, 42] to Llama. We ask the LLM to generate summary counterfactuals for missing key steps [K-CF] in which some crucial action steps (clip narrations) are missing,
@@ 3.2 Preliminary: Hierarchical Video
In this work, we adopt HierVL [2], a framework that learns hierarchical video-language representations at two temporal scales, clip-level and video-level. The clip-level narrations and video-level summaries are annotated by Ego4D/Egoclip [23, 42]. Note that since HierVL trains both a visual and text encoder, it requires vision-to-text and text-to-vision alignment. Below we describe the former, and the latter is defined symmetrically. Formally, let a video be divided into $M$ short clips {$v_{1},\ldots,v_{M}$} with corresponding narration texts {$t_{1},\ldots,t_{M}$}, and let $S$ be the video-level summary for the entire video. HierVL’s vision-to-text training objective consists of two parts: clip-level alignment and video-level alignment losses.
@@
where $B$ is the mini-batch, $v_{i}$ and $t_{j}$ are visual and textual embeddings, respectively, and $P(i)$ denotes the positive samples of the $i^{\text{th}}$ video clip. HierVL [2] leverages a variant of the action-aware loss, EgoNCE [42], in order to mine hard positives and negatives; please refer to [2, 42] for more details about this sampling. The overall loss at the child level is thus $\mathcal{L}_{\text{child}}=\mathcal{L}_{{v2t}}+\mathcal{L}_{{t2v}}$, where the latter is defined symmetrically.
@@
Br-prompt [39] GTEA 94.1 92.0 83.0 91.6 81.2 88.4 - - - -
I3D [7] Kinetics 90.1 88.8 79.2 84.6 79.7 84.5 48.8 71.9 73.9 64.9
CLIP [56] WIT [75]+Text encoder 88.5 86.2 77.6 87.1 75.6 83.0 44.2 71.2 70.8 62.1
MIL-NCE [50] HowTo100M 67.9 61.3 44.6 67.9 58.3 60.0 47.3 69.1 73.6 63.3
PVRL [103] HowTo100M 85.2 82.6 72.2 81.1 71.2 78.5 45.6 73.2 73.4 64.1
HierVL [2] Ego4D+Text encoder 90.4 88.5 81.2 86.7 78.5 85.1 52.6 73.0 77.3 67.6
@@ 3.3 Pretraining Objective:
Frame-Level Alignment. At the frame level, the model is supervised by our proposed before-state loss $\mathcal{L}_{\text{before}}$ and after-state loss $\mathcal{L}_{\text{after}}$, enhancing action-induced transformation for procedural understanding. The mathematical formulation of each is [29],
@@
Clip-Level Alignment. At the clip-level, we leverage the $L_{v2t}$ loss described in Section 3.2, which seeks to align the video-clip embeddings to their corresponding text narrations from EgoClip [42]. Note that since we do not train a text encoder, the symmetric $L_{t2v}$ loss is neglected here. For more details on this loss, see Section 3.2 and [2, 42]. The resulting loss for the first two scales is therefore
@@
where $V_{i}$ is the aggregated video-level visual embedding, $S_{j}$ is a summary text embedding and $S_{j,w}^{cf}$ are Misordered and Missing-step counterfactual text embeddings where $w\in\{1,\ldots,W\}$ is the total number of counterfactuals used, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ video, respectively. We also use a temperature hyperparameter $\tau$ as in Eq. (3), but we omit it in the equation above for simplicity of notation. Note that consistent with [2], in this case, the summation over positive samples is located inside the logarithm since the sampling strategy of [42] is used. However, prior work [29] finds the summation outside the logarithm to be more effective and thus this is the de-facto choice of the frame-level loss in Eq. (3).